OpenStudy (anonymous):

Guys need help to get the answer for the below qn? please help.... A particle is moving along the curve given by the equation r(t) = 2t i + t^2 j +lnt k a) find velocity of the particle b) the speed c) distance travelled from t=0 d) acceleration e) tangential acceleration f) the acceleration normal to the tangent and g) the radius of curvature any hope????

OpenStudy (sirm3d):

velocity is \(r'(t)\) speed = \(|r'(t)|\) distance traveled =\(\displaystyle \int_0^{t} |r'(s)|\mathrm ds\)

OpenStudy (aravindg):

hope for the best :) ...welcome to openstduy !

OpenStudy (sirm3d):

acceleration is \(r''(t)\)

OpenStudy (sirm3d):

tangential acceleration = \(\displaystyle \frac{r''(t)\cdot r'(t)}{|r'(t)|^2}r'(t)\)

OpenStudy (sirm3d):

acceleration normal = acceleration - tangential acceleration

OpenStudy (sirm3d):

radius of curvature \(\rho(t)\) \[\Large\rho(t)=\frac{|r'(t)|^3}{|r'(t)\times r''(t)|}\] have fun. =)

OpenStudy (anonymous):

great thanks